# GEX: Democratizing Dexterity with Fully-Actuated Dexterous Hand and Exoskeleton Glove

### v1

@@ III-B Dexterous Retargeting
Due to the structural inconsistency between the dexterous hand and the exoskeleton glove, dexterous retargeting is required during teleoperation to accurately map the EX12-worn human hand movements onto the GX11’s joint space. To address this, we developed a retargeting system based on Dex-Retargeting [^3] https://github.com/dexsuite/dex-retargeting, with the source code being open-sourced [^4] https://github.com/Democratizing-Dexterous/gex_retargeting_sim. The retargeting results in the PyBullet simulation environment are illustrated in Fig. 7. As evident from the figure, the dexterous retargeting algorithm achieves precise mapping of pinching operations, which is crucial for dexterous teleoperation tasks.

### v2

@@ III-B Dexterous Retargeting
Due to the structural inconsistency between the dexterous hand and the exoskeleton glove, dexterous retargeting is required during teleoperation to accurately map the EX12-worn human hand movements onto the GX11’s joint space. To address this, we developed a retargeting system based on Dex-Retargeting [^3] https://github.com/dexsuite/dex-retargeting, with the source code being open-sourced [^4] https://github.com/Democratizing-Dexterous/gex_retargeting_sim.
The retargeting is framed as constrained, with the objective of finding a suitable joint position trajectory $\textbf{q}^{rbt}_{t}$ that minimizes the following cost function:
$\min_{\mathbf{q}^{rbt}_{t}}\sum_{i=1}^{n_{f}}\|F^{rbt}_{i}(\mathbf{q}^{rbt}_{t})-\alpha F^{exo}_{i}(\mathbf{q}_{t}^{exo})\|^{2}+k\|\mathbf{q}_{t}^{rbt}-\mathbf{q}_{t-1}^{rbt}\|^{2}$ (1)
$\text{s.t.}\quad q_{l}\leq q^{rbt}_{t}\leq q_{u},$
where the first summation term enforces spatial pose consistency by minimising the squared deviation between the fingertip pose of the $i^{\text{th}}$ finger of the dexterous hand, $F^{rbt}_{i}(\mathbf{q}^{rbt}_{t})$, and the corresponding fingertip pose of the exoskeleton, $F^{exo}_{i}(\mathbf{q}^{exo}_{t})$; here $n_{f}$ denotes the number of force channels and $\alpha\in\mathbb{R}^{+}$ is a dimension-matching scaling factor. The second term $k\lVert\mathbf{q}^{rbt}_{t}-\mathbf{q}^{rbt}_{t-1}\rVert^{2}$ penalises inter-step variations in joint positions with a smoothing weight $k>0$, thereby suppressing trajectory jitter and enhancing executability. The bound constraint $q_{l}\leq q^{rbt}_{t}\leq q_{u}$ (with $q_{l},q_{u}\in\mathbb{R}^{n_{q}}$ denoting joint lower and upper limits, respectively) explicitly guarantees adherence to the robot’s mechanical and safety limits. The keypoint placement of the exoskeleton and the dexterous hand is shown in Fig. 7, where the keypoints are located at the fingertips and at the base of the palm. The retargeting results of the exoskeleton under different hand poses are illustrated in Fig. 9. As shown in the figure, the exoskeleton motions in various configurations are consistently mapped to the corresponding postures of the dexterous hand, demonstrating the effectiveness of the retargeting scheme.
Figure 7: Task space vectors between fingertips and palm for both the teleoperation glove and GX11 pro hand used for retargeting optimization.
The retargeting results in the PyBullet simulation environment are illustrated in Fig. 8. As evident from the figure, the dexterous retargeting algorithm achieves precise mapping of pinching operations, which is crucial for dexterous teleoperation tasks.
@@
Figure 9: Visualization of retargeting results in Isaac from glove to dexterous hand. The first row shows the glove poses, and the second row shows the retargeted results on the dexterous hand. The green spheres in the glove represent the positions of the human fingertips.
@@ IV Experiments
Figure 12: Illustration of the three manipulation tasks using GX11 hand: grasping, reorientation, and twisting. The top row shows table-top grasping, where the robot picks up a target object from the table and moves it to a goal region. The middle row shows object reorientation on the table. The bottom row shows in-hand twisting of an object mounted on a stand. Our dexterous hand is able to successfully accomplish all three tasks within a unified control framework.
@@
IV-A Manipulation tasks
As illustrated in Fig. 12, we evaluate our method on three representative dexterous manipulation tasks: grasping, reorientation, and twisting. In the grasping task, the robot arm and dexterous hand must approach a target object on the table, establish a stable grasp, and transport the object to a designated goal area. In the reorientation task, the hand interacts with a composite object lying on the tabletop and changes its global pose from a random initial configuration to a desired orientation through a sequence of contact-rich motions. In the twisting task, the hand grasps an object mounted on a stand and performs a controlled rotational motion around the specified axis while maintaining a stable grasp. Across these three tasks, our dexterous hand successfully completes the required behaviors, demonstrating its ability to handle both pick-and-place style operations and fine in-hand manipulation.